OpenStudy (anonymous):

How many square inches of a material is needed to create an open top rectangular solid whose volume is 54 cubic inches? Assume that the length and width are of the same measurement and that the height is twice the length.

OpenStudy (unklerhaukus):

\[w=l=\frac h 2\qquad\qquad V=wlh\qquad\qquad S=2(wl+wh+lh)\]

OpenStudy (unklerhaukus):

w=width, l=lenght, h=height, V=volume, S=surface area

OpenStudy (anonymous):

thank you :D

OpenStudy (unklerhaukus):

actually it is an open top so\[S=(wh+lh)\]

OpenStudy (anonymous):

let 54-volume x-length x-width 2x-height 54= L x W x H 54= (X)(X)(2X) 54=2x^3 x=3 distribute all of them : )

OpenStudy (anonymous):

S=((3)(3)+(3)(6)) S=9+18 S=27